# Exploring the evolution of red and blue galaxies in different cosmic web environments using IllustrisTNG simulation

### v1

@@ 4.8 Evolution of sSFR across diverse cosmic web environments
The median sSFR across different environments is shown in Figure 17. The top left panel indicates that, unlike the median SFR, there are no statistically significant differences in the median sSFR across environments in the redshift range $31.5<z<$ 3. As redshift decreases, sSFR decreases in all environments. At $z<1$, clusters exhibit the lowest median sSFR, with filaments showing somewhat higher sSFR compared to clusters between $z=1$ and $z=0$. Sheets and voids have the highest and similar median sSFR in this redshift range. The top right panel presents results for the lower mass bin, which mirrors the trends observed in the entire sample for the reasons discussed earlier. The bottom panel of Figure 17 displays results for the higher mass bin. This panel shows similar trends but with larger differences in median sSFR across environments. It is interesting to note that the median sSFR for massive galaxies in voids tends to increase between $z=2$ and $z=1$, though this trend is difficult to confirm due to large errorbars for the galaxies in voids.
@@ 5 Discussions and Conclusions
(i) We track the evolution of the red and blue fractions across different cosmic web environments from redshift $z=3$ to $z=0$. Initially, blue galaxies are more prevalent in clusters compared to other geometric environments. This trend reverses at lower redshifts. The red fraction starts to rise after specific redshifts for each environment, with clusters experiencing earlier suppression of star formation compared to filaments, sheets, and voids. The massive galaxies show a consistent increase in red fractions over time in all environments.
(ii) We analyze the red and blue fractions as a function of stellar mass across different environments at various redshifts. At $z=3$, clusters have a higher blue fraction and lower red fraction compared to other environments at all masses, with filaments showing an intermediate blue fraction. Over time, the red fraction increases in higher mass galaxies and decreases in lower mass galaxies at all environments until $z=1.5$. A pronounced environment-dependent trend emerges at $z<1$ during which all environments show a significant increase in the red fraction at higher masses. On the other hand, clusters show a significant rise whereas other environments display a moderate increase in the red fraction at lower masses. The quenching in lower mass galaxies is more influenced by their geometric environment compared to the higher mass galaxies.
(iii) We study the relative fractions of red and blue galaxies across different cosmic web environments. Filaments and sheets show a decrease in the relative red fraction (RRF) and an increase in the relative blue fraction (RBF) with decreasing redshift, while clusters exhibit the opposite trend. For lower mass galaxies, the trends in RRF and RBF are consistent with the entire sample. For higher mass galaxies, clusters initially have higher RRF and RBF than filaments. This trend reverses around $z\sim 1$. The differences in RRF and RBF across different geometric environments are primarily decided by the varying proportions of galaxies in these environments.
(iv) We investigate how the RRF and RBF of galaxies change with stellar mass and redshift in various cosmic web environments. Initially, the RRF is higher for lower mass galaxies in filaments compared to those in clusters. However, as redshift decreases, the RRF in clusters gradually surpasses that in filaments. The RRF in clusters at lower masses increases with decreasing redshift, making clusters dominant for low-mass red galaxies and filaments for intermediate masses. The RRF for clusters dominates at higher masses at all redshifts. It does not change significantly across the redshift range $3-0$, implying that massive ($\log(\frac{M_{*}}{M_{\odot}})>11$) red galaxies are mostly found in galaxy clusters. At lower masses, the RRF in clusters remains between $\sim(30-40)\%$ in this redshift range. On the other hand, RBF in filaments consistently have a higher relative proportion ($\sim(50-60)\%$) of blue galaxies, particularly at lower masses. This suggests that they host a diverse galaxy population, most likely due to less intense environmental quenching. The clusters, despite being densest, exhibit highest RBF at higher masses, indicating potential rejuvenation of star formation through interactions or mergers.
(v) The $(u-r)$ colour distribution of galaxies is unimodal at higher redshift and gradually shifts towards redder values over time, eventually showing bimodality by $z=2$, with the red peak growing most in clusters. At $z<1$, clusters have the highest median colour followed by filaments, sheets, and voids. The median colours of massive galaxies do not differ significantly across various environments in this redshift range implying that colour evolution is largely driven by mass rather than environment.
(vi) Clusters exhibit highest median mass followed by filaments, sheets and voids at all redshifts. The median stellar mass of galaxies increases with decreasing redshift across all environments. We observe a slower growth of median mass for lower mass galaxies in clusters at $z<1$. In contrast, the median mass for higher mass galaxies remain stable during $z=3-1.5$ and gradually increases at $z<1$ in all environments. The median SFR declines with decreasing redshift in all environments, with clusters dominating until $z=1.5$. At $z<1.5$, the median SFR in clusters levels out across environments and then decreases relative to others. Notably, the median SFR for massive galaxies drops to zero in all environments at $z<1.5$, indicating suppression of star formation in high-mass galaxies. The median sSFR decreases in all environments with decreasing redshifts with no significant differences across environments until $z=1.5$. At $z<1.5$, clusters exhibit the lowest median sSFR and voids and sheets exhibiting the highest values. The higher mass galaxies show larger variations in sSFR across environments.
Our study suggests that both stellar mass and geometric environments play an important role in galaxy evolution. The higher density regions host a larger fraction of blue galaxies during $z=3-2$, implying they provide a more favourable environment for star formation. However, the environmental factors start to dominate at $z<2$ reversing this trend. We find that at $z<1$, most of the massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) are quenched irrespective of their geometric environments. This is consistent with AGN feedback model in the IllustrisTNG simulations which suggests that AGN feedback is highly effective in massive central galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$) [128]. Kinetic feedback from super-massive black holes (SMBHs) is the key mechanism driving the quenching of star formation in massive galaxies [129]. The fraction of quenched galaxies in TNG300 is in the range $(30-80)\%$ percent in the stellar mass range $10^{10.5}-10^{11.5}M_{\odot}$ [127]. Our results (Figure 3, Figure 4) are consistent with these findings. So the AGN feedback may have played a dominant role in quenching the high mass galaxies in all environments. However, the AGN feedback model in IllustrisTNG shows reduced effectiveness in quenching galaxies of intermediate and lower masses [128]. At lower masses ($\log(\frac{M_{*}}{M_{\odot}})<10.5$), quenching is largely controlled by environment. The fraction of red galaxies at lower masses increases with decreasing redshift at each geometric environment. At lower masses, clusters host the highest fraction of red galaxies followed by filaments, sheets and voids. Low mass galaxies often undergo environmental quenching before they become part of their current host group. The pre-processing can involve interactions with other galaxies or the effects of the surrounding medium, leading to a reduction in their gas reservoirs and star formation rates.
The relative fractions of red and blue galaxies (RRF and RBF) are influenced by both stellar mass and geometric environment. Generally, RRF increases and RBF decreases in clusters as redshift decreases. In contrast, filaments and sheets exhibit the opposite trend. At higher masses, both RRF and RBF in clusters decline with decreasing redshift, while in filaments, both show a steady increase. This anomalous behavior can be largely attributed to a growing relative abundance of intermediate and high-mass galaxies in filaments. Interestingly, clusters, despite being the densest environments, display the highest RRF and RBF for galaxies with masses $\log(\frac{M_{*}}{M_{\odot}})>11$. This suggests that massive galaxies are predominantly located in clusters, where most are quenched, although some may be rejuvenated through interactions or mergers. In contrast, filaments host a more diverse galaxy population as evidenced by the higher RBF and RRF at lower masses ($\log(\frac{M_{*}}{M_{\odot}}<11$). At lower masses, filaments exhibit relative red fractions (RRF) comparable to clusters, yet host nearly $60\%$ of blue galaxies. This suggests that the milder environmental quenching in filaments enables galaxies to undergo a wider range of evolutionary stages. A recent study [90] using the SIMBA simulation [91] suggest that shock heating in filaments can inhibit star formation. Another study [92] with the IllustrisTNG simulation finds that satellite galaxies are especially susceptible to such quenching. Our results are consistent with these findings. The low-mass quenched galaxies in filaments may represent satellite populations, while the star-forming galaxies may correspond to central galaxies with intermediate masses.
Our study reveals several key trends in galaxy evolution across different geometric environments over time. At higher redshifts, the $(u-r)$ colour distribution is unimodal but turns into bimodal by $z=2$, particularly in clusters, which have the highest median colour at $z<1$. The median colour and stellar mass of galaxies increases steadily from $z=3$ ato $z=0$. For massive galaxies, a steady increase in median colour across all environments, with little variations across environments, suggests that the process driving the reddening of massive galaxies is fairly uniform across the cosmic web. This could indicate that the quenching mechanisms, such as AGN feedback, which cause galaxies to evolve from blue to red, operate similarly across different environments. In contrast, the significant differences in median stellar mass across environments suggest that the growth of galaxy mass is more sensitive to the local cosmic web environment, with different environments supporting different rates of stellar mass buildup. This might reflect environmental factors influencing galaxy mergers or accretion more strongly than colour evolution. This is further supported by the decline in SFR and sSFR across all environments, with massive galaxies showing a sharp reduction in these rates at lower redshifts, regardless of their environment.
We also compare our results with the earlier studies on galaxy evolution with hydrodynamical simulations. [109] investigate the evolution of galaxy colour using the EAGLE simulation [130] and find that the red sequence forms around $z\sim 1$ due to the quenching of low-mass satellite galaxies and AGN feedback in more massive central galaxies. [111] demonstrate using TNG100 and TNG300 that between $70\%$ and $90\%$ of central and satellite galaxies with masses greater than or equal to $\log(\frac{M_{*}}{M_{\odot}})>10$ are predicted to be quenched in the redshift range $0\leq z\leq 0.5$. [110] demonstrate that in the EAGLE simulation, the quenching time scales for both central and satellite galaxies decrease with increasing stellar mass. They observe an increase in the fraction of passive galaxies at lower redshifts. Another study [112] using the TNG simulation finds that slow-quenching galaxies are approximately twice as common as fast-quenching galaxies in the redshift range of $20.7<z<$ 2. More recently, [113] study the roles of local density and interactions on the evolution of red and blue galaxies in the redshift range $0\leq z\leq 3$ using the EAGLE simulation. They report that red galaxies tend to occupy denser regions and show significantly stronger clustering compared to blue galaxies at $z<2$. They also find a substantial decrease in the blue fraction in paired galaxies at $z<1$ suggesting significant roles of interactions in galaxy evolution. Our results are consistent with these findings. Further, Our findings also align with the observational studies [99, 97, 131, 98], reinforcing the effectiveness of hydrodynamical simulations as a valuable tool for investigating galaxy formation and evolution.
Our study indicates that quenching at higher masses is mainly driven by stellar mass, whereas quenching in intermediate- and low-mass galaxies is influenced by a complex interplay between stellar mass and environment. We conclude that both stellar mass and cosmic web environments play a significant role in shaping colour bimodality and driving overall galaxy evolution.
@@ ACKNOWLEDGEMENTS
BP would like to acknowledge IUCAA, Pune, for providing support through the associateship programme. AN acknowledges the financial support from the Department of Science and Technology (DST), Government of India through an INSPIRE fellowship. AN thanks Dylan Nelson for help in downloading and understanding the data from the IllustrisTNG database.

### v2

@@ 4.8 Evolution of sSFR across diverse cosmic web environments
The median sSFR across different environments is shown in Figure 18. The top left panel indicates that, unlike the median SFR, there are no statistically significant differences in the median sSFR across environments in the redshift range $1.5<z<3$. As redshift decreases, sSFR decreases in all environments. At $z<1$, clusters exhibit the lowest median sSFR, with filaments showing somewhat higher sSFR compared to clusters between $z=1$ and $z=0$. Sheets and voids have the highest and similar median sSFR in this redshift range. The top right panel presents results for the lower mass bin, which mirrors the trends observed in the entire sample for the reasons discussed earlier. The bottom panel of Figure 18 displays results for the higher mass bin. This panel shows similar trends but with larger differences in median sSFR across environments. It is interesting to note that the median sSFR for massive galaxies in voids tends to increase between $z=2$ and $z=1$, though this trend is difficult to confirm due to large errorbars for the galaxies in voids.
@@ 5 Discussions and Conclusions
(i) The evolution of red and blue fractions show strong environmental dependence. At $z=3$, clusters host the highest fraction of blue galaxies, but this trend reverses at lower redshifts as environmental quenching becomes dominant. The inflection points at $z=1.5$ for clusters, $z=0.7$ for filaments and sheets, and $z=0.5$ for voids mark key transitions where red galaxies start to dominate in these environments.
(ii) The environmental quenching plays a major role at low masses. For low-mass galaxies ($\log(\frac{M_{*}}{M_{\odot}})<10.5$), clusters host the highest red fractions at $z<1$, showing that environmental effects drive quenching in these systems. Galaxies in filaments exhibit intermediate quenching signatures, spanning a broader range of evolutionary stages. At $z<1$, low-mass galaxies in filaments show a higher blue fraction compared to those in clusters.
(iii) The filaments serve as a bridge between the galaxy evolution in cluster and void type environments. At $z<1$, filaments host a mix of quiescent and star-forming galaxies, with nearly $60\%$ of low-mass galaxies remaining blue. The presence of mixed galaxy populations in filaments may reflect both the less extreme environmental conditions, potentially allowing more gradual quenching, and the fact that filaments host a large fraction of the galaxy population, thus capturing a broad diversity of evolutionary stages.
(iv) Our findings support the interpretation that stellar mass plays a significant role in the quenching of high-mass galaxies. For massive galaxies ($\log(\frac{M_{*}}{M_{\odot}})>10.5$), quenching occurs independent of environment, suggesting that internal processes like mass quenching and AGN feedback dominate over external effects.
(v) The sharp increase in the red fraction at $\log(\frac{M_{*}}{M_{\odot}})\sim 10.5$ marks a transition in galaxy evolution. It corresponds to a buildup of intermediate-mass galaxies and represents a shift from star-forming to quiescent populations, becoming more prominent at $z<1$.
(vi) The colour distribution transitions from unimodal ($z>2$) to bimodal ($z<2$), with clusters developing the strongest red peak. The strengthening of colour bimodality over time indicates the build-up of quiescent galaxies at lower redshifts.
(vii) The median colours of massive galaxies do not differ significantly across various environments implying that colour evolution in high mass galaxies is largely driven by mass rather than environment. The median mass of lower mass galaxies in clusters exhibit a slower growth at $z<1$. This indicates that environmental factors influence galaxy mergers or accretion more strongly than colour evolution.
Our findings align well with previous observational and simulation-based studies on galaxy evolution within the cosmic web, each probing specific redshift regimes and scales. For example, studies such as [133] (SDSS, $0.02\leq z\leq 0.155$), [134] (VIPERS, $z\sim 0.7$), [135] (COSMOS, $0.5<z<0.9$), [136] (GAMA, $0.03\leq z\leq 0.25$), [137] (WISExSCOS, $0.1<z<0.3$), [138] (SAMI, $0.004<z<0.13$), and [139] (WISE x SCOSMOS, $0.1<z<0.3$) report that galaxies near filament spines tend to be redder, more massive, and less star-forming at fixed local density on scales of a few to several Mpc. However, as pointed out by [140] (SDSS, $z\leq 0.077$), some of these trends may be fully explained by local density effects alone, without invoking additional influence from the filament geometry.
One widely accepted interpretation is that galaxies migrating toward filament spines experience elevated accretion rates, more frequent interactions, and increased merger activity leading to accelerated mass growth and quenching. This interpretation is supported by simulation-based work across various redshifts and scales, including IllustrisTNG and zoom-in hydrodynamical studies [87, 141, 142, 143, 144, 145, 146]. Additional processes, such as pre-processing in small galaxy groups embedded within filaments [147, 148] (z=0.150.7z=0.15-0.7), and potential shock heating in filaments suppressing star formation especially in satellite galaxies are also implicated by recent analyses using SIMBA and IllustrisTNG simulations [149, 150] (z=02z=0-2). Our observation of an elevated relative red fraction in filaments at intermediate stellar masses may be connected to these combined mechanisms.
We also reinforce the picture of filaments as transitional environments. Observational studies such as [151] and [152], spanning $0.5\lesssim z\lesssim 2$, show that low-mass galaxies in filaments maintain higher star formation rates than their cluster counterparts at fixed stellar mass and local density. High-resolution zoom-in simulations by [153] find that anisotropic gas accretion onto filament halos at high redshifts ($z=2.5,4$) enhances their baryon and stellar fractions relative to field analogs. Dark matter-only and hydrodynamical simulations [154, 124] (z=03z=0-3) similarly indicate that low-mass halos preferentially accrete perpendicularly to their host filament, while higher-mass halos accrete along the filament axis. Observations of HI content in SIMBA simulations [155] (z=06z=0-6) and related anisotropic inflow signatures [156] (z=02z=0-2) further reinforce this scenario.
Finally, the transition in colour distribution from unimodal at $z>2$ to bimodal at $z<1$ supports findings from the EAGLE simulation [105, 106], which identified buildup of the red sequence around $z\sim 1$ due to AGN feedback and satellite quenching. The increasing red fractions at $\log(M_{*}/M_{\odot})>10.5$ are also consistent with the mass-dependent quenching trends reported in [130, 157, 131]. Overall, our results align with the broader consensus that environment-driven quenching dominates for low-mass galaxies, while mass- and AGN-driven quenching become increasingly important at higher masses [128, 129]. Throughout our analysis, the cosmic web classification is based on a Gaussian-smoothed galaxy density field with a smoothing scale of 4 Mpc. This choice is designed to probe large-scale structures such as filaments and sheets, which typically span several megaparsecs in width. We emphasize that our results pertain to this intermediate scale and may not be directly comparable to studies that focus on much larger-scale structures (e.g., superclusters) or smaller-scale environments (e.g., galaxy groups or cluster cores). Care should be taken when interpreting similarities or differences across studies that operate at significantly different spatial resolutions or employ alternative definitions of cosmic web environments.
The cosmic web serves as the primary source of fuel for galaxy growth, with its filamentary network playing a crucial role in galaxy formation and evolution. Our findings emphasize the intricate interplay between galaxy mass and environment, both of which collectively shape the evolutionary pathways of galaxies within the cosmic web.
@@ ACKNOWLEDGEMENTS
Authors thank an anonymous reviewer for insightful comments and suggestions that helped us to improve the draft. BP would like to acknowledge IUCAA, Pune, for providing support through the associateship programme. AN acknowledges the financial support from the Department of Science and Technology (DST), Government of India through an INSPIRE fellowship. AN thanks Dylan Nelson for help in downloading and understanding the data from the IllustrisTNG database.
@@
Appendix A Appendix
We examine how local density and filter width affect our results, with detailed findings for different density bins and filter widths presented in the following two subsections.
A.1 The impact of varying local density on galaxy fractions
Figure 19: This shows the redshift evolution of total fraction (F), red fraction (RF), blue fraction (BF), relative red fraction (RRF), and relative blue fraction (RBF) across different cosmic web environments. Each row corresponds to one specific fraction, while each column shows results for a narrow bin in local density, $\rho$, with density ranges labeled at the top.
We would like to disentangle the relative influence of local density and cosmic web environment on galaxy colour evolution. We examine the redshift evolution of total fraction (F), red fraction (RF), blue fraction (BF), relative red fraction (RRF), and relative blue fraction (RBF) in different cosmic web environments across 5 narrow local density bins in Figure 19. This allows a controlled comparison of these fractions at fixed densities. Each row presents the redshift evolution of one fraction, while each column corresponds to a specific density range. Differences observed across environments within the same density bin reveal the distinct impact of the geometric environments of the cosmic web, independent of local density.
We note that within each local density bin, clear and consistent differences persist between the cosmic web environments which indicates that cosmic web geometry influences galaxy evolution independently of local density. In the top row of Figure 19, which shows the total fraction (F) of galaxies in each cosmic web environment, a consistent pattern emerges across all density bins. In each local density bin, filaments consistently host the largest fraction of galaxies, followed by sheets, with voids and clusters contributing smaller fractions. As we move from low to high density bins (left to right), the fraction of galaxies in clusters gradually increases, while the fraction in sheet decreases, as expected. However, filaments dominate across all bins, regardless of local density. This suggests that filaments are the most common environment for galaxies across the full range of densities examined. The fraction of galaxies in clusters dominates that in filaments only at higher densities and higher redshifts.
The second row of Figure 19 shows that RF increases over time across all environments, with the rate of increase depending on both local density and cosmic web geometry. In all density bins, clusters consistently exhibit the highest red fractions, followed by filaments, sheets, and voids, especially at lower redshifts. As local density increases, overall RF values rise, but the ordering of environments remains stable, indicating that geometric environment influences quenching beyond what local density alone accounts for. An exactly opposite trend is observed in the RF at higher redshifts where RF is highest in voids followed by sheets, filaments and clusters. These results imply that both local density and cosmic web environment shape the build-up of red galaxies, with geometry modulating quenching efficiency even at fixed density. The third row of the Figure 19 shows the BF, which is complementary to the RF and reflects the relative abundance of star-forming galaxies in each environment and density bin. These results are directly related to the results presented in the second row of Figure 19 and support the interpretation that star formation is more sustained in underdense environments at lower redshift, while it is more efficient in denser and clustered regions at higher redshifts.
Figure 20: This shows the impact of Gaussian filter width on the redshift evolution of total fraction (F), red fraction (RF), blue fraction (BF), relative red fraction (RRF), and relative blue fraction (RBF) across different cosmic web environments. Each row represents one galaxy fraction, while each column corresponds to a different filter width.
The fourth (RRF) and fifth (RBF) rows of Figure 19 reveal trends that closely mirror those in the total fraction (F) shown in the first row, and this similarity provides deeper insight into the interplay between galaxy population demographics and environmental influence within fixed local density bins. In all density bins, filaments consistently dominate the total galaxy population, as shown in the first row. Correspondingly, in both the RRF and RBF, filaments also contribute the largest relative share of both red and blue galaxies, particularly at intermediate densities and redshifts. This co-dominance arises not necessarily from an intrinsic tendency for filaments to favour one galaxy type, but from the sheer number of galaxies they host, positioning them as the primary channel for both ongoing star formation and quenching within the cosmic web. The alignment between total fraction (F), RRF and RBF highlights an important point. Geometric environment shapes not just galaxy properties, but also where galaxies reside, and these distributions, in turn, amplify the environment’s role in galaxy evolution. Filaments are central to this narrative. They act as both the main reservoirs of galaxies and the transitional sites where galaxies experience a mix of quenching and star formation. Clusters and voids represent the extremes, favouring quenching and sustained star formation respectively. In essence, the combined trends in RRF, RBF, and F underscore that galaxy colour evolution is not merely about local density or intrinsic processes, but is profoundly modulated by where galaxies live in the cosmic web. Filaments, due to their prevalence and mixed influence, play a pivotal and nuanced role in shaping galaxy populations. The shifting cross-overs in the first, fourth, and fifth rows show that the influence of cosmic web environments on galaxy properties evolves with local density. At low densities, filaments dominate in both total fraction and relative red and blue fractions, but as density increases, clusters overtake filaments in hosting red galaxies, reflecting enhanced environmental quenching. These trends highlight a dynamic interplay between local density and geometric environment, where density modulates but does not replace the impact of cosmic web environment.
Figure 19 clearly demonstrates that galaxy colour evolution is not solely governed by local density, but is also significantly influenced by the cosmic web environment. Across all density bins, distinct differences persist in the red and blue fractions among voids, sheets, filaments, and clusters. These differences cannot be explained by density alone. Our results suggest that galaxy colour evolution is driven by both local density and cosmic web environment. While local density controls the overall quenching trend, the geometric environment modulates the efficiency and timing of quenching, even at fixed local density.
A.2 The impact of varying filter width on galaxy fractions
We study the influence of the finite width of the Gaussian filter on our results in this subsection. The Figure 20 shows that the filter width used in cosmic web classification significantly influences the observed galaxy fractions and their evolutionary trends. While the overall qualitative behaviors persist across different filter widths, where clusters dominate red fractions, voids dominate blue fractions, and filaments host the majority of galaxies. However the relative contributions shift with filter width. Smaller filter widths better resolve transitional environments like filaments, crucial for understanding gradual quenching processes. At smaller widths (2 Mpc), the structures in the cosmic web are more finely resolved, allowing filaments and sheets to play a prominent role in hosting both red and blue galaxies. In contrast, a larger filter width smooth over finer structures blurring the boundaries between high-density peaks and surrounding structures. As the filter width increases, clusters contribute less to the RRF and RBF. This indicates that clusters become less distinct in the smoothed density field, likely being diluted into surrounding filaments and sheets due to loss of high-resolution contrast. Filaments and sheets, in contrast, gain a larger share of both red and blue galaxies at larger filter widths These results highlight that filter width is not a neutral choice. It modulates the apparent dominance of different environments and thus shapes our interpretation of how galaxies evolve within the cosmic web. Therefore, the choice of filter width should be carefully tailored to the characteristics and resolution of the dataset being analyzed.
Based on these findings, we adopt a Gaussian smoothing scale of 4 Mpc for our main analysis. This choice reflects a careful balance between spatial resolution and structural coherence. A 4 Mpc filter is slightly below the mean intergalactic separation in our sample ($\sim 5$ Mpc), allowing us to resolve filamentary and sheet-like structures without introducing excessive fragmentation or noise. At this scale, transitional environments such as filaments are sufficiently preserved, enabling a meaningful analysis of gradual quenching and mixed galaxy populations. In contrast, smaller smoothing scales (e.g., 2 Mpc) introduce finer detail but can amplify shot noise, while larger scales tend to over-smooth dense structures, diluting cluster cores into their surroundings. Moreover, a 4 Mpc filter aligns with the typical width of filaments observed in both simulations and large-scale galaxy surveys, ensuring that our results remain physically interpretable and directly comparable to previous studies. Overall, this smoothing length provides an optimal trade-off for capturing the complexity of the cosmic web in a statistically robust and physically meaningful way.